OpenStudy (anonymous):

If y=√x\lnx, then y′(e^2)=? use quotient rule

OpenStudy (anonymous):

HI :DDD

OpenStudy (anonymous):

\[\large y=\frac{\sqrt{x}}{\ln(x)}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\left(\frac{f}{g}\right)'=\frac{gf'-fg'}{g^2}\] with \[f(x)=\sqrt{x},f'(x)=\frac{1}{2\sqrt{x}},g(x)=\ln(x), g'(x)=\frac{1}{x}\] most of it will be annoying algebra

OpenStudy (anonymous):

first step is just to plug that stuff in to the quotient rule, start with \[\large\frac{\ln(x)\times\frac{1}{2\sqrt{x}}-\sqrt{x}\times \frac{1}{x}}{(\ln(x))^2}\]

OpenStudy (anonymous):

got it, yea i was a bit confused on how to set it up. thanks!

OpenStudy (anonymous):

next step is probably to get rid of obnoxious compound fraction by multiplying top and bottom by \(2x\)

OpenStudy (anonymous):

yw